# Primordial Black Hole Interpretation in Subsolar Mass Gravitational Wave Candidate SSM200308

### v2

@@ II Bayesian analysis of SSM200308
$\mathcal{R}_{12,3}(t\mid\!\vec{\theta})$ $\approx\frac{7.9\times 10^{4}}{\mathrm{Gpc}^{3}\mathrm{yr}}\left(\frac{t}{t_{0}}\right)^{\frac{\gamma}{7}-1}f_{\mathrm{PBH}}^{\frac{144\gamma}{\frac{259}{}}+\frac{47}{37}}$ (2)
$\times$ ${\left[\frac{\langle m\rangle}{M_{\odot}}\right]^{\frac{5\gamma-32}{37}}\left(\frac{M}{2\langle m\rangle}\right)^{\frac{1799}{259}-\frac{2122}{333}}(4\eta)^{-\frac{3\gamma}{7}-1}}$
@@ III.2 SGWB form binary PBH coalescences
Figure 3: The SIGW (purple lines) and SGWB from binary PBHs (blue lines) if SSM200803 can be explained by PBHs. The violin polot represents the NANOGrav 15yr dataset Agazie et al. (2023). The purple and blue bands represent the the region at $95\%$ confidence level for SIGWs and SGWB from binary PBHs respectively.

### v3

@@ II Bayesian analysis of SSM200308
$\mathcal{R}_{12,3}(t\mid\!\vec{\theta})$ $\approx\frac{7.9\times 10^{4}}{\mathrm{Gpc}^{3}\mathrm{yr}}\left(\frac{t}{t_{0}}\right)^{\frac{\gamma}{7}-1}f_{\mathrm{PBH}}^{\frac{144\gamma}{259}+\frac{47}{37}}$ (2)
$\times$ ${\left[\frac{\langle m\rangle}{M_{\odot}}\right]^{\frac{5\gamma-32}{37}}\left(\frac{M}{2\langle m\rangle}\right)^{\frac{179\gamma}{259}-\frac{2122}{333}}(4\eta)^{-\frac{3\gamma}{7}-1}}$
@@ III.2 SGWB form binary PBH coalescences
Figure 3: The SIGW (purple lines) and SGWB from binary PBHs (blue lines) if SSM200803 can be explained by PBHs. The violin polot represents the NANOGrav 15yr dataset Agazie et al. (2023). The purple and blue bands represent the region at $95\%$ confidence level for SIGWs and SGWB from binary PBHs respectively.